Bubbly:

Which best describes the triangle, if any, that can be formed with sides measuring 60 centimeters, 100 centimeters, and 40 centimeters?

Bubbly:

A.two unique scalene triangles B.infinetly many similar scalene triangles C.no triangle D.one unique scalene triangle

Bubbly:

@kittybasil

barstho:

Hm, I'm not really good with trigonometry. But we know that a scalene triangle is a triangle where all 3 sides have different lengths. I'm inclining towards D, but I'm not completely sure. Take someones else opinion over mine if you get it. (:

Bubbly:

Thats okay, thank you. I'll take all the help I can get.

barstho:

Let's break it down a bit

Bubbly:

okay

barstho:

C is definitely out of the question

kittybasil:

Hm? I got tagged in this. Actually, C isn't necessarily out of the question

barstho:

@kittybasil Hm, how so?

kittybasil:

Well, what is your reasoning behind the idea that it must make a triangle? 🤔 Technically, if it does not meet the Triangle Inequality Theorem then it cannot form a triangular shape.

kittybasil:

Triangle Inequality theorem explains that only certain lengths that fit into the formula can become sides of a triangle.

barstho:

Right. So the sum of any two sides must be greater than the measure of the 3rd side (off of google)

Bubbly:

So using this theorem would it be A then

barstho:

60 + 40 = 100 Which means it's not greater than the 3rd side. Shouldn't that mean it can't form a triangle?

Bubbly:

but 60+100=160 and 40+100=140??

Bubbly:

so two triangles?

kittybasil:

Basically, you replace the variables in the formula with any of the number values given to you. If at least one combination works, then that combination will give you a triangle.

kittybasil:

Your values are: 60 cm, 100 cm, and 40 cm. Is there any way you can make the equation work with these numbers?

Bubbly:

6 by using the theorem?

Bubbly:

because when I used the pythagorean theorem it doesn't form a triangle.

kittybasil:

Pythagorean theorem is only for right angle triangles. Anything else requires the Triangle Inequality theorem

Bubbly:

oh so then it would be a because when I did it 60+100=160 and 40+100=140

kittybasil:

Let's make the variables equal to certain values to avoid confusion: A = 40 B = 100 C = 60 Based on this, if ANY of these following expressions are true using the values above, then that expression makes a triangle. A + B > C B + C > A A + C > B

Bubbly:

the last one is equal to b so does that mean the answer is c?

barstho:

Knew it all along 😌 @kittybasil

Bubbly:

but I thought it said if one wasn't greater than one side there is no triangle

kittybasil:

You are correct

Bubbly:

who?

kittybasil:

you are. 😂

Bubbly:

oh, okay, thank you guys for your help!

barstho:

No problem, I think i confused you guys more than helped, but I learned something new. (:

kittybasil:

Making an effort is the first step in learning. You guys all did something, but I can't medal everyone :p